evaluate each expression.
This is the problem I need help with: ((( 5(a - b) for a= -1/2 and b= -2/3)))

Answer by edjones(8007)

5(a-b)
=5a-5b
=-5/2 - (-10/3)
=-5/2 + 10/3
=-15/6 + 20/6
=5/6
.
Ed